OpenStudy (anonymous):

Show that the point on a unit circle associated to the angle t=pie/3 is (1/2, the square root of 3/2)

OpenStudy (jdoe0001):

http://upload.wikimedia.org/wikipedia/commons/4/45/30-60-90_triangle.jpg by 30-30-60 rule, the "hypotenuse" is twice the length of the "adjacent" side the rules also points out that the "opposite" side would be "adjacent times root 3" but once you have the adjacent and the hypotenuse, you can just use the pythagorean theorem to get the opposite side $$ c^2 = a^2+\color{red}{b}^2 \implies \color{red}{b}^2 = \sqrt{c^2-a^2} $$

OpenStudy (jdoe0001):

\(\large \cfrac{\pi}{2} = 60^o\)

OpenStudy (jdoe0001):

darn I meant \(\large \cfrac{\pi}{3} = 60^o\)

OpenStudy (anonymous):

pie/2 is 90 degrees i thought our teacher asked up to prove this with the distance formula.

OpenStudy (jdoe0001):

my bad typo :/

OpenStudy (anonymous):

its all good!

OpenStudy (anonymous):

would you know how to make it work with the distance formula...? I have figure it out some the part that i an struggling with would be the (x-(the square root of 3/2) square part.. :/

OpenStudy (jdoe0001):

the distance formula is a "renamed" version of the pythagorean theorem, they both do exactly the same, different name :), so use the distance formula, just don't call it pythagorean

OpenStudy (anonymous):

\[\sqrt{(x-0)^{2}+(y-1)^{2}}=\sqrt{(x-\sqrt{3}/2)^{2}+(y-1/2)^{2}}\]

OpenStudy (jdoe0001):

well, for the distance formula you'd just need to have to points to work on, for \(\frac{\pi}{3} \) you'd use the origin as one (0,0) and some other point, the endpoint for the \(\frac{\pi}{3}\) angle

OpenStudy (jdoe0001):

$$ d = \sqrt{(x_2-x_1)+(y_\color{blue}{2}-y_1)}\\ d^2 = (x_2-x_1)+(y_\color{blue}{2}-y_1)\\ \text{now using the origin and 1/2 for }x_2\\ \text{keeping in mind the distance between then is "1"}\\ 1^2 = \cfrac{1}{4}+y_\color{blue}{2}^2 \implies 1 -\cfrac{1}{4} = y_2^2\\ y_\color{blue}{2}^2=\sqrt{\cfrac{3}{4}} \implies y_\color{blue}{2}^2=\cfrac{\sqrt{3}}{\sqrt{4}} = \cfrac{\sqrt{3}}{2} $$ that is, using a 2nd point of \(\pmatrix{\frac{1}{2}, y_2}\)

OpenStudy (jdoe0001):

hmmm, yet another typo :/

OpenStudy (jdoe0001):

$$ d = \sqrt{(x_2-x_1)+(y_\color{blue}{2}-y_1)}\\ d^2 = (x_2-x_1)+(y_\color{blue}{2}-y_1)\\ \text{now using the origin and 1/2 for }x_2\\ \text{keeping in mind the distance between then is "1"}\\ 1^2 = \cfrac{1}{4}+y_\color{blue}{2}^2 \implies 1 -\cfrac{1}{4} = y_2^2\\ y_\color{blue}{2}=\sqrt{\cfrac{3}{4}} \implies y_\color{blue}{2}=\cfrac{\sqrt{3}}{\sqrt{4}} = \cfrac{\sqrt{3}}{2} $$

OpenStudy (jdoe0001):

there